# Thermalization in Krylov Basis

### v1

@@ Abstract
We study thermalization for closed non-integrable quantum systems in the Krylov basis. Following the idea of the eigenstate thermalization hypothesis, one may introduce Krylov basis thermalization hypothesis which imposes a condition on the matrix elements of local operators in the Krylov basis. In this context, the nature of thermalization may be probed by the infinite time average of the Krylov complexity. We also compute the variance of Lanczos coefficients which may provide another quantity to examine the nature of thermalization. We will see that there is a direct relation between the behavior of the infinite time average of complexity and that of the inverse participation ratio of initial states.
@@ I Introduction
Even though for closed quantum systems one may also observe emerging of the thermal equilibrium in non-equilibrated systems (caused by e.g. global quench), unlike the classical systems, the thermalization may be seen without performing any time averages Srednicki:1994mfb ; Deutsch:1991 . Indeed, out of equilibrium states approach to their thermal expectations short after relaxation. It is, however, important to note that in closed quantum systems dynamics is unitary and time reversal invariant, and therefore, a priori, it is not obvious how and in what sense the thermal equilibrium can be reached dynamically.
@@
Although, it is believed that a non-integrable model will generally thermalize, the nature of thermalization might be different in different situations. Actually, besides Hamiltonian which gives dynamics of the system, the nature of the thermalization may also depend on the initial state, such that, within a fixed model different initial states may exhibit different behaviors Banuls:2011vuw .
@@
which suggests that the information of the effective inverse temperature could be read from the expectation value of the energy. Indeed, the regime on which the strong or weak thermalization may occur could be also identified by the normalized energy of the initial state Chen:2021
@@
where $|Z\pm\rangle$ are eigenvectors of $\sigma^{z}$ with eigenvalues $\pm$. Indeed, at each site the corresponding state is the eigenvector of the operator ${\cal O}_{i}=n\cdot\sigma_{i}$, with $n$ is the unit vector on the Bloch sphere. More explicitly, one has
@@
Using this analytic expression for the expectation value of energy we have drown the density of energy in figure 1 for $N=100$ and $J=1,\,h=0.5,\,g=-1.05$. Actually, to highlight the regions where the density of energy vanishes we have depicted its absolute value.
@@
One observes that the behavior of the effective inverse temperature matches exactly that of the density of energy event though the size of the two systems by which these quantities are evaluated are different by about a factor of 15. This shows the robustness of the results against the size of the system. In particular, this has to be compared with the results in the literature where the numerical computations have been performed for $N=14$. Even though our $\beta$ is evaluated for $N=7$ in comparison with that of $N=14$ the error we acquire is about 3 percent.
The aim of the present article, which will be explored in the next section, is to study quantum thermalization using the Krylov basis which seems more appropriate basis when the dynamics of the system is of our interest. Within this context, we will also study the nature of thermalization.
@@ II Krylov basis and thermalization
In the energy eigenstates the expectation value (11) reads
$\langle{\cal O}(t)\rangle=\overline{\langle{\cal O}\rangle}+\sum_{n\neq m}^{\cal D}e^{i(E_{n}-E_{m})t}c_{n}c^{*}_{m}\langle E_{m}|{\cal O}|E_{n}\rangle\,$ (12)
where ${\cal D}$ is the dimension of Hilbert space, $c_{n}$ is coefficient of expansion of the initial state in energy basis and bar stands for infinite time average: $\overline{\langle{\cal O}\rangle}=\frac{1}{T}\int_{0}^{T}\langle{\cal O}(t)\rangle$ for $T\rightarrow\infty$ limit. Then, one can proceed to explore equilibrium and thermalization in this context which happens due to the possible phase cancellation in long times Serdnicki:1999 .
@@
II.1 Thermalization in Krylov basis
Here we would like to study quantum thermalization in the Krylov basis. In particular, we would like to understand to what extent the nature of thermalization may be explored in this context. To proceed, let us first briefly review the recursive procedure producing the Krylov space for a given state in a quantum system ( see viswanath1994recursion for review).
@@ II.1 Thermalization in Krylov basis
This recursive procedure stops whenever $b_{n}$ vanishes which occurs for $n={\cal D}_{\psi}\leq{\cal D}$ that is the dimension of Krylov space. Recall that ${\cal D}$ is the dimension of the Hilbert space of the quantum system. Note that this procedure produces an orthonormal and ordered basis together with coefficients $a_{n}$ and $b_{n}$ known as the Lanczos coefficients Lanczos:1950zz .
@@
The infinite time average of the corresponding operator is given by
$\overline{\langle{\cal O}\rangle}=\sum_{m,n=0}^{{\cal D}_{\psi}-1}C_{nm}O_{nm}\,,$ (18)
where
@@
is a complex matrix with the conditions $C_{nm}^{*}=C_{mn}$ and $\sum_{n}C_{nn}=1$. Following the idea of ETH, in our framework, in order to get an infinite time average consistent with the macro canonical result, one may assume that the matrix elements, $O_{nm}$, should satisfy an ETH like condition
$O_{nm}\sim O\delta_{mn}+{\cal O}({\cal D}^{-1}),$ (20)
by which we are guaranteed to get the expected thermal equilibrium for the expectation value of the corresponding operator. Here $O$ is a constant. One may want to call this remarkable behavior as Krylov basis thermalization hypothesis (KTH).
@@
To examine KTH behavior of local operators (observables) we will compute the matrix elements of the operator $S_{x}=\sum_{i=1}^{N}\sigma^{x}_{i}$, which is magnetization in the $x$ direction, for the Ising model (1) with $h=0.5,\,g=-1.05$ where the model is non-integrable.
@@
The corresponding matrix elements for two different initial states $|Y+\rangle$ and $|Z+\rangle$ are presented in figure 3 for $N=10$. From this figure, one observes that the matrix elements $O_{nm}$ exhibit desired behavior as imposed by KTH indicating that thermalization occurs for both cases, as expected.
@@
Figure 3: Matrix elements of the operator $S_{x}=\sum_{i}\sigma_{i}^{x}$ in Krylov basis for cases where the initial state is $|Y+\rangle$ (left) and $|Z+\rangle$ (right). Darker points represent the matrix elements that are closed to zero.
To highlight the behavior of matrix elements we have presented the absolute value of them, so that in figure 3 the dark points correspond to vanishing elements. Note that, in general, the off diagonal matrix elements are complex numbers and what has been depicted in this figure is the real part of the matrix elements.
In comparison with ETH where matrix elements of operators in energy eigenstates are real, having complex off diagonal elements in KTH makes the situation more involved. Indeed, one should also examine the imaginary part of the off diagonal matrix elements that contribute to off diagonal terms in the expectation value of the operators. It also requires to study of the behavior of $C_{nm}$ as well.
@@
We have also computed matrix elements of the magnetization in the $z$ direction, $S_{z}=\sum_{i=1}^{N}\sigma^{z}_{i}$, for several initial states specified by different $\theta$ and $\phi$ and we have found the same pattern as that in figure 3. Although in all cases we have considered, our numerical computations support the KTH proposal, we think it is interesting to explore this point better work .
@@
To further explore the KTH behavior we have also done the same computations for the cases where $gh=0$ in which the model (1) is integrable. An immediate observation we have made is that in integrable cases the dimension of Krylov space reduces significantly [^4] Similar observation has been already made in the context of operator Krylov complexity in Rabinovici:2020ryf (see also Caputa:2024vrn ). which may also depend on the initial state. For example, for $N=10$ while in the chaotic case the dimension of Krylov space is 529, in an integrable mode given by $h=0,\;g=-1.05$ it is 463 for initial states $|Y+\rangle,\,|Z+\rangle$ and 253 for $|X+\rangle$. For integrable model given by $h=0.5,\;g=0$ it is 23 for $|Y+\rangle,\,|X+\rangle$ and zero for $|Z+\rangle$.
@@
In order to have a better statistic we have considered the case where $h=0,\,g=-1.05$. For this case, we have computed matrix elements of different operators in the Krylov basis for different initial states. We have found that, although for some special cases the corresponding matrix elements have almost similar pattern as that in figure 3, it is not a generic behavior and typically they exhibit non-universal behavior. Actually, even in some cases the real part of the matrix elements are identically zero.
It is worth noting that there is a special operator in Krylov space whose matrix elements is exactly diagonal. More precisely, consider the number operator defined by
${\cal N}=\sum_{n=0}^{{\cal D}_{\psi}-1}\,n|n\rangle\langle n|,$ (21)
that is obviously diagonal in Krylov basis, ${\cal N}_{nm}=n\delta_{nm}$. In this case the expectation value (11), actually, computes Krylov complexity Parker:2018yvk
that saturates (equilibrate) at very late times where the Lanczos coefficients vanish Rabinovici:2022beu .
II.2 Krylov space and nature of thermalization
Restricting ourselves to the non-integrable phase, as it is clear from the above construction, the Lanczos coefficients should contain information about both the dynamics of the model and the initial state making it a suitable candidate to probe the nature of thermalization. Therefore, it is natural to see if they could also probe weak and strong thermalization. In particular, although both initial states $|Y+\rangle$ and $|Z+\rangle$ exhibit almost the same pattern for the operator matrix elements, indicating that both states yield to thermalization, we are wondering if it is possible to distinguish between the nature of thermalization for these states using the properties of Lanczos coefficients.

### v3

@@ Abstract
We study thermalization in closed non-integrable quantum systems using the Krylov basis. We demonstrate that for thermalization to occur, the matrix representation of typical local operators in the Krylov basis should exhibit a specific tridiagonal form with all other elements in the matrix are exponentially small, reminiscent of the eigenstate thermalization hypothesis. Within this framework, we propose that the nature of thermalization, whether weak or strong, can be examined by the infinite time average of the Krylov complexity. Moreover, we analyze the variance of Lanczos coefficients as another probe for the nature of thermalization. One observes that although the variance of Lanczos coefficients may capture certain features of thermalization, it is not as effective as the infinite time average of complexity.
@@ I Introduction
Even though for closed quantum systems one may also observe emerging of the thermal equilibrium in non-equilibrated systems (caused by e.g. global quench), unlike the classical systems, the thermalization may be seen without performing any time averages Srednicki:1994mfb ; Deutsch:1991 . Indeed, out of equilibrium states approach to their thermal expectations shortly after relaxation. It is, however, important to note that in closed quantum systems dynamics is unitary and time reversal invariant, and therefore, a priori, it is not obvious how and in what sense the thermal equilibrium can be reached dynamically.
@@
Although, it is believed that a non-integrable model will generally thermalize, the nature of thermalization might differ in different situations. Actually, besides Hamiltonian which gives dynamics of the system, the nature of the thermalization may also depend on the initial state, such that, within a fixed model different initial states may exhibit different behaviors Banuls:2011vuw .
@@
which suggests that the information of the effective inverse temperature could be read from the expectation value of the energy. Indeed, the regime on which the strong or weak thermalization may occur could also be identified by the normalized energy of the initial state Chen:2021
@@
where $|Z\pm\rangle$ are eigenvectors of $\sigma^{z}$ with eigenvalues $\pm$. Indeed, at each site, the corresponding state is the eigenvector of the operator ${\cal O}_{i}=n\cdot\sigma_{i}$, with $n$ is the unit vector on the Bloch sphere. More explicitly, one has
@@
Using this analytic expression for the expectation value of energy we have drawn the density of energy in figure 1 for $N=100$ and $J=1,\,h=0.5,\,g=-1.05$. Actually, to highlight the regions where the density of energy vanishes we have depicted its absolute value.
@@
One observes that the behavior of the effective inverse temperature matches exactly that of the density of energy event though the size of the two systems by which these quantities are evaluated are different by about a factor of 15. This shows the robustness of the results against the size of the system. In particular, this has to be compared with the results in the literature where the numerical computations have been performed for $N=14$. Even though our $\beta$ is evaluated for $N=7$ in comparison with that of $N=14$ the error we acquire is about ${\cal O}(1)$ percent.
This article aims to study quantum thermalization using the Krylov basis which seems to be a more appropriate basis when the dynamics of the system is our interest. Although the Krylov method has been used to study numerical computations viswanath1994recursion , in recent years there have been several activities to use the Krylov method in the context of quantum chaos (see Parker:2018yvk and its citations).
In this paper, we would like to explore a potential application of Krylov space within the context of quantum thermalization. The key advantage of studying thermalization in this framework lies in the fact that, under a unitary time evolution the trajectory of a given initial state does not necessarily expose into the entire Hilbert space. Instead, it remains confined within a subset known as the Krylov space, which typically has a smaller dimension compared to the full Hilbert space of the system. Thus, focusing on the Krylov space suffices for studying the time evolution of the system. In particular, if the system has conserved charges, working in this basis we are automatically confined in a subsystem that preserves the symmetry of the initial state.
The paper is organized as follows. In the next section, we will study the late time behavior of the expectation value of typical operators within the Krylov basis. Following the Eigenstate Thermalization Hypothesis (ETH), we will propose an ansatz for the matrix elements of the operator within the Krylov basis, called the Krylov Thermalization Hypothesis (KTH). We will present several numerical computations in support of our ansatz. In section three we will study the nature of thermalization in this framework. Specifically, we will introduce two metrics to probe the nature of thermalization—namely, the variance of Lanczos coefficients and the infinite time average of Krylov complexity. While the former may not provide a definitive conclusion, the latter offers a pattern that agrees perfectly with other proposed probes in the existing literature. Additionally, we will calculate the inverse participation ratio to contrast it with the results from the infinite time average of complexity. The late section is devoted to discussions.
@@ II Krylov basis and thermalization
In the energy eigenstates, assuming $|\psi_{0}\rangle=\sum_{\alpha}c_{\alpha}|E_{\alpha}\rangle$, the expectation value (11) reads
$\langle{\cal O}(t)\rangle={\rm Tr}(\rho_{DE}{\cal O})+\sum_{\alpha\neq\beta}^{\cal D}e^{i(E_{\alpha}-E_{\beta})t}c_{\alpha}c^{*}_{\beta}\langle E_{\alpha}|{\cal O}|E_{\beta}\rangle\,,$ (12)
where $\rho_{DE}$ is diagonal density matrix
$\rho_{DE}=\sum_{\alpha=1}^{\cal{D}}|c_{\alpha}|^{2}|E_{\alpha}\rangle\langle E_{\alpha}|\,.$ (13)
Here ${\cal D}$ is the dimension of Hilbert space. Then, one can proceed to explore equilibrium and thermalization in this context which happens due to the possible phase cancellation at long times Serdnicki:1999 when the expectation value of the operator may be given by the canonical ensemble ${\rm Tr}(\rho_{DE}{\cal O})\approx{\rm Tr}(\rho_{th}{\cal O})$.
@@
Here we would like to study quantum thermalization in the Krylov basis. In particular, we would like to understand to what extent the nature of thermalization may be explored in this context (see also Bhattacharjee:2022qjw ). To proceed, let us first briefly review the recursive procedure producing the Krylov space for a given state in a quantum system ( see viswanath1994recursion for review).
@@
This recursive procedure stops whenever $b_{n}$ vanishes which occurs for $n={\cal D}_{\psi}\leq{\cal D}$ that is the dimension of Krylov space. Note that this procedure produces an orthonormal and ordered basis together with coefficients $a_{n}$ and $b_{n}$ known as Lanczos coefficients Lanczos:1950zz .
@@
Using the completeness of the energy eigenstates one may expand any element of the Krylov basis in terms of energy basis
$|n\rangle=\sum_{\alpha=1}^{{\cal D}}f_{n\alpha}|E_{\alpha}\rangle\,.$ (18)
Note that since ${\cal D}_{\psi}\leq{\cal D}$, the expansion coefficient, $f_{n\alpha}$, is not necessary inevitable and therefore, in general, energy eigenstates cannot be expanded in terms of Krylov basis. Using the orthogonality condition of the Krylov basis one gets
$\sum_{\alpha=1}^{{\cal D}}f^{*}_{n\alpha}f_{m\alpha}=\delta_{nm}\,.$ (19)
On the other hand from the equation (14) one finds
$f_{n\alpha}E_{\alpha}=a_{n}f_{n\alpha}+b_{n+1}f_{n+1\alpha}+b_{n}f_{n-1\alpha}\,,$ (20)
that can be used to find $f_{n\alpha}$ in terms of $f_{0\alpha}=c_{\alpha}$.
@@
To find the infinite time average of the corresponding operator one needs to compute $C_{nm}$ given by
@@
that is essentially matrix elements of the diagonal density matrix in the Krylov basis $C_{nm}=\langle m|\rho_{DE}|n\rangle$, that is
$C_{nm}=\sum_{\alpha=1}^{\cal{D}}|c_{\alpha}|^{2}f_{m\alpha}f^{*}_{n\alpha}\,.$ (23)
In this framework, inspired by ETH, we propose an ansatz for the matrix elements of typical operators in the Krylov basis to ensure that thermalization occurs within the system. To illustrate how the corresponding ansatz might be formulated, let us consider the ETH ansatz for the matrix elements of the operator ${\cal O}$ in energy eigenstates Srednicki:1994mfb ; Serdnicki:1999 444 In this expression $S(\bar{E})$ is thermal entropy at energy $\bar{E}$ which is an extensive quantity and proportional to the size of the system. It is important to note that $f$ and $f_{O}$ are smooth functions of their arguments. $R_{\alpha\beta}$ is a random real or complex variable with zero mean $\overline{R_{\alpha\beta}}=0$ and unit variance: $\overline{R^{2}_{\alpha\beta}}=1,\,\overline{|R_{\alpha\beta}|^{2}}=1$.,
$\langle E_{\alpha}|{\cal O}|E_{\beta}\rangle=f(E_{\alpha})\delta_{\alpha\beta}+e^{-\frac{S(\bar{E})}{2}}f_{{\cal O}}(\bar{E},\omega)R_{\alpha\beta}\,,$ (24)
where $\bar{E}=(E_{\alpha}+E_{\beta})/2,\;\omega=E_{\alpha}-E_{\beta}$. This expression can be utilized to propose an ansatz for the matrix elements of the operator ${\cal O}$ in the Krylov basis through the following relation
$\langle n|{\cal O}|m\rangle=\sum_{\alpha,\beta=1}^{{\cal D}}f^{*}_{n\alpha}f_{m\beta}\;\;\langle E_{\alpha}|{\cal O}|E_{\beta}\rangle\,.$ (25)
To proceed, one can promote the function $f$ to an operator by replacing the energy with the Hamiltonian
$f(E_{\alpha})\rightarrow\hat{f}(H)\,.$ (26)
This allows to express the ETH ansatz as follows
$\langle E_{\alpha}|{\cal O}|E_{\beta}\rangle=\langle E_{\alpha}|\hat{f}(H)|E_{\beta}\rangle+{\cal O}(e^{-S/2})\,.$ (27)
Here we have utilized the fact that $\hat{f}(H)|E_{\beta}\rangle=f(E_{\beta})|E_{\beta}\rangle$. Plugging this expression into equation (25) one gets
$\langle n|{\cal O}|m\rangle=\langle n|\hat{f}(H)|m\rangle+{\rm suppressed\;terms}\,.$ (28)
In this equation, the suppressed terms correspond to the exponentially suppressed terms in the original ETH ansatz. Using the fact that $\langle n|m\rangle=\delta_{nm}$ and
$\langle n|H|m\rangle=a_{n}\delta_{n\;m}+b_{m+1}\delta_{n\;m+1}+b_{m}\delta_{n\;m-1},$ (29)
it becomes clear that the matrix representation of the operator ${\cal O}$ in the Krylov basis is not diagonal. However, we note that while off-diagonal elements also appear in leading order in this expression, we generally would not expect significant contributions from all off-diagonal matrix elements.
To understand this, we recognize that for thermalization to occur beside the ETH ansatz, one must further assume that $f$ is a smooth and slowly varying function of $E_{\alpha}$. Additionally, the initial state must be sufficiently localized within a narrow energy window—specifically, the variance of energy should be much smaller than the energy expectation value of the initial state. This justifies the neglecting of higher-order terms in the following Taylor expansionSerdnicki:1999  [^5] More precisely, one assumes that $(\Delta E)^{2}f^{\prime\prime}(E)/f(E)\ll 1$ with $\Delta E$ being the variance of energySerdnicki:1999 .
$f(E_{\alpha})\approx f(E)+(E_{\alpha}-E)\,f^{\prime}(E),$ (30)
where $E=\langle\psi_{0}|H|\psi_{0}\rangle$ and, $f(E)$ is the expectation value predicted by the (micro)canonical ensemble ${\rm Tr}(\rho_{th}{\cal O})\approx f(E)$. Here “prime” denotes derivative with respect to $E_{\alpha}$. In this approximation, the ETH ansatz (24) reads
${\cal O}_{\alpha\beta}\approx f(E)\delta_{\alpha\beta}+f^{\prime}(E)\,\langle E_{\alpha}|H-E|E_{\beta}\rangle+{\cal O}(e^{-\frac{S}{2}})\,,$ (31)
resulting in the following expression for the matrix elements in the Krylov basis
${\cal O}_{nm}$$\approx$ $f(a_{0})\delta_{nm}+f^{\prime}(a_{0})\,\langle n|H-a_{0}|m\rangle$ (33)
$+{\rm suppressed\;terms}\,,$
which shows that the matrix representation of a typical observable in the Krylov basis is essentially tridiagonal in which the off-diagonal elements adjacent to the diagonal, denoted as ${\cal O}_{nn+1}$, are proportional to the Lanczos coefficients $b_{n}$ (see equation (29)). Note that the suppressed terms, which contain $b_{n}$-dependent factors, are of order ${\cal O}(\Delta E^{2})$.
Regarding ”suppressed terms,” one can follow a similar approach as with the ETH ansatz to estimate their order of magnitude DAlessio:2015qtq . To proceed, following Mori:2017qhg , we note that using the completeness of the Krylov basis, the inequality the fact $\langle n|{\cal O}^{2}|n\rangle\leq|{\cal O}|^{2}$ maybe written as $\sum_{m}|{\cal O}_{nm}|^{2}\leq|{\cal O}|^{2}$. Moreover, since $|{\cal O}_{nn}|^{2}$ is always positive one may write
$\sum_{m(\neq n)=0}^{{\cal D}_{\psi}-1}|{\cal O}_{nm}|^{2}\leq|{\cal O}|^{2}\,.$ (34)
where $|{\cal O}|$ denotes the operator norm666 The norm is defined by $|{\cal O}|={\rm sup}_{\psi}\sqrt{\langle\psi|{\cal O}^{\dagger}{\cal O}|\psi\rangle}$ Mori:2017qhg . , and ${\cal D}_{\psi}$ is the dimension of the Krylov space associated with the initial state $\psi_{0}$. Moreover, if we assume that the off-diagonal matrix elements ${\cal O}_{nm}$ are smooth and vary slowly, they can be considered to be almost constant so that one finds
$\sum_{m(\neq n)=0}^{{\cal D}_{\psi}-1}|O_{nm}|^{2}\approx|O_{nm}|^{2}\sum_{m(\neq n)=0}^{{\cal D}_{\psi}-1}=|O_{nm}|^{2}{\cal D}_{\psi}\,.$ (35)
By making use of this approximation one can derive an expression that captures the behavior of off-diagonal elements to the leading order. Indeed from equation (34), we one gets
$|{\cal O}_{nm}|\leq\frac{|{\cal O}|}{\sqrt{{\cal D}_{\psi}}}\,.$ (36)
It is essential to highlight that the aforementioned condition pertains to the contributions of ”suppressed terms” to off-diagonal matrix elements. Additionally, there is a contribution to these off-diagonal matrix elements from leading terms, as indicated in equation (33).
Inspired by the above observations one can propose an ansatz for the matrix elements of typical operators in the Krylov basis as follows
$\langle n|{\cal O}|m\rangle=\langle n|\hat{f}(H)|m\rangle+\frac{1}{\sqrt{{\cal D}_{\psi}}}f_{{\cal O}}(a_{nm})R_{nm}\,,$ (37)
where $a_{nm}=\langle n|H|m\rangle$. Here $R_{nm}$ is a random real or complex variable with zero mean $\overline{R_{nm}}=0$ and unit variance: $\overline{R^{2}_{nm}}=1,\,\overline{|R_{nm}|^{2}}=1$.
The equation (37) imposes a condition on the matrix elements of the operator ${\cal O}$ in the Krylov basis, which can be seen as an ansatz for these matrix elements to ensure thermalization. This is referred to as the KTH ansatz. It is worth noting that, in practice, for typical chaotic systems, the leading-order terms of the KTH ansatz are actually represented by those in equation (33). In fact by substituting this expression into equation (21) one finds
$\langle{\cal O}(t)\rangle=\sum_{n,m=0}^{{\cal D}_{\psi}-1}\phi^{*}_{n}(t)\phi_{m}(t)\,\langle n|\hat{f}(H)|m\rangle$ (38)
$\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{1}{\sqrt{{\cal D}_{\psi}}}\sum_{n,m=0}^{{\cal D}_{\psi}-1}\phi^{*}_{n}(t)\phi_{m}(t)f_{{\cal O}}(a_{nm})R_{nm}\,.$
Note that this equation indicates that the off-diagonal terms, which contain $b_{n}$-dependent terms, remain exponentially suppressed, scaling as ${\cal O}(e^{-S/2})$.
One can use (33) to simplify the first line of the above equation. It is clear that the contribution of off-diagonal term in (33) vanishes [^7] Note that $\sum_{n,m=0}^{{\cal D}_{\psi}-1}\phi^{*}_{n}(t)\phi_{m}(t)\,\langle n|H|m\rangle=E$ and by definition $a_{0}=E$., whereas from the first term and taking into account that $\sum_{n}|\phi(t)|^{2}=1$ one arrives at
$\langle{\cal O}(t)\rangle\approx f(a_{0})+\frac{1}{\sqrt{{\cal D}_{\psi}}}\sum_{n,m=0}^{{\cal D}_{\psi}-1}\phi^{*}_{n}(t)\phi_{m}(t)\,f_{{\cal O}}(a_{nm})R_{nm},$ (39)
which results in
$\langle{\cal O}(t)\rangle\approx{\rm Tr}(\rho_{th}{\cal O})+{\rm small\;fluctuations},$ (40)
as expected. Here we have used that the system thermalizes so that the expectation value of the operator is equal to thermal expectation value $f(a_{0})={\rm Tr}(\rho_{th}{\cal O})$. Therefore, even though the matrix is tridiagonal, the main contribution is primarily determined by the diagonal elements, which represent the expectation value of the corresponding operator as derived from the (micro)canonical ensemble. It is also worth noting that in our context, all physical results are fundamentally determined by two parameters: $a_{0}$ and $b_{1}$, both of which carry physical significance. Specifically, $a_{0}=E$ represents the energy of the initial state, while $b_{1}=\Delta E$ denotes its variance..
@@
To examine the KTH behavior of local operators (observables) we will compute matrix elements of the operator $S_{x}=\sum_{i=1}^{N}\sigma^{x}_{i}$, which is magnetization in the $x$ direction, for the Ising model (1) with $h=0.5,\,g=-1.05$ where the model is non-integrable.
@@
The corresponding matrix elements for two different initial states $|Y+\rangle$ and $|Z+\rangle$ are presented in figure 3 for $N=10$. From this figure, one observes that the matrix elements $O_{nm}$ exhibit the desired behavior as suggested by KTH. To highlight the behavior of matrix elements we have presented the absolute value of them, so that in figure 3 the dark points correspond to vanishing elements.
@@
Figure 3: Matrix elements of the operator $S_{x}=\sum_{i}\sigma_{i}^{x}$ in Krylov basis for cases where the initial state is $|Y+\rangle$ (left) and $|Z+\rangle$ (right). Darker points represent the matrix elements that are closed to zero. Although, due to its resolution, it might not be clear for these plots, these matrices are not diagonal and, indeed, they are tridiagonal (see figure 4).
To further explore KTH we have also presented the actual values of matrix elements of $S_{x}$ in the Krylov basis for different initial states in figure 4 where one can observe that $(S_{x})_{n,n+1}$ is significantly greater than other elements. Note also that, diagonal elements ${\cal O}_{nn}$ are not entirely given by $f(a_{0})$ and, in fact, they appear in a certain combination of $f(a_{0})+(a_{n}-a_{0})f^{\prime}(a_{0})$ which might be small, even though the expectation value, itself, could be relatively large.
Figure 4: Actual (absolute) values of matrix elements of the operator $S_{x}=\sum_{i}\sigma_{i}^{x}$ in Krylov basis for cases where the initial state is $|Y+\rangle$ (up) and $|Z+\rangle$ (down). Blue points in left panels denote $(S_{x})_{n,n+1}$ elements.
@@
We have also computed matrix elements of the magnetization in the $z$ direction, $S_{z}=\sum_{i=1}^{N}\sigma^{z}_{i}$, for several initial states specified by different $\theta$ and $\phi$ and we have found the same pattern as that in figure 3 and 4. In particular, numerical results for initial states $|Y+\rangle$ and $|Z+\rangle$ are depicted in figure5.
Figure 5: Actual (absolute) values of matrix elements of the operator $S_{z}=\sum_{i}\sigma_{i}^{z}$ in Krylov basis for cases where the initial state is $|Y+\rangle$ (up) and $|Z+\rangle$ (down). Blue points in the left panels denote $(S_{z})_{n,n+1}$ elements.
It is also illustrative to explicitly compute time evolution of the operator we considered above to see how they actually follow the general behavior given by the equation (40). The results are depicted in figure 6. The straight brown lines in these plots represent the long time expectation value predicted by the canonical ensemble which is equal to the infinite time average of the corresponding expectation value: ${\rm Tr}(\rho_{DE}S_{x,z})\approx{\rm Tr}(\rho_{th}S_{x,z})$, that is necessary for thermalization to occur.
Figure 6: Time evolution of the expectation value of $S_{x}$ and $S_{z}$ for two different initial states $|Y+\rangle$ (up) and $|Z+\rangle$ (down). The straight lines represent the value predicted by the canonical ensemble.
Although thermalization occurs for both initial states, from this figure one observes that the nature of thermalization should be different for these states. While $|Y+\rangle$ exhibit string thermalization, for $|Z+\rangle$ it is weak.
@@
To further explore the KTH behavior we have also done the same computations for the cases where $gh=0$ in which the model (1) is integrable. An immediate observation we have made is that in integrable cases the dimension of Krylov space reduces significantly [^8] Similar observation has been already made in the context of operator Krylov complexity in Rabinovici:2020ryf (see also Caputa:2024vrn ). which may also depend on the initial state. For example, for $N=10$, where the dimension of the full Hilbert space is $2^{10}$, in the chaotic case, the dimension of the Krylov space is 529999It is known that the Hamiltonian (1) has a parity symmetry which is essentially reflection symmetry about the center of the chain. It is straightforward to see that the initial state (6) has positive parity which in turn indicates that the obtained Krylov subspace should be a subspace with positive parity. Actually, the positive parity subspace has 528 dimensions, as expected, which is equal to the dimension of Krylov space.. In the integrable case with parameters $h=0$ and $g=-1.05$, the corresponding dimensions are 463 for the initial states $|Y+\rangle$ and $|Z+\rangle$, and 253 for $|X+\rangle$.
@@
In order to have a better statistic we have considered the case where $h=0,\,g=-1.05$. For this case, we have computed matrix elements of different operators in the Krylov basis for different initial states. We have found that, although for some special cases, the corresponding matrix elements have almost similar patterns as that in figure 3, it is not a generic behavior and typically they exhibit non-universal behavior. More importantly, in this case the long time average cannot be approximated by a canonical ensemble (see, for example, figure 7).
Figure 7: Time evolution of the expectation value of $S_{x}$ for initial state $|Y+\rangle$ for an integrable case where $h=0$. The brown dashed line represents a value predicted by the canonical ensemble whose energy is the same as that of the initial state $|Y+\rangle$.
III Krylov space and nature of thermalization
As we have already maintained although both initial states $|Y+\rangle$ and $|Z+\rangle$ exhibit almost the same pattern for the operator matrix elements (see figures 4 and 5) indicating that thermalization occurs in both states, it is evident for figure 6 that the nature of thermalization for these two states must be different, as we discussed in the previous section. In this section, we would like to study how the nature of thermalization, being weak or strong, can be probed in the context of Krylov space.
From equation (40) one finds that the nature of thermalization should be controlled by the second term which is essentially given by a summation over $\phi_{n}(t)$’s. On the other hand, $\phi_{n}(t)$ can be evaluated, recursively, from $\phi_{0}(t)$ using Lanczos coefficients. Therefore, the nature of thermalization should be reflected in these quantities. Based on this insight, we will propose various quantities within the framework of Krylov space that could serve as indicators for the nature of thermalization.
III.1 Variance of Lanczos coefficients
From the Krylov basis construction, it is evident that the Lanczos coefficients should encapsulate information about both the model’s dynamics and initial state, making them a suitable candidate for probing the nature of thermalization.